OpenStudy (hellomynameiskam):

Which ratio proportional to 8/18 A. 4/5 B. 16/26 C. 2/5 D. 20/15

OpenStudy (denonakavro):

Divide both 8 and 18 by 2 8/2 = 4 18/2 = 9 or you can multiply them both by 2 8 x 2 =16 18 x 2 = 36 So the answer should either be 4/9 or 16/36 So there is a mistake with your choices

OpenStudy (hellomynameiskam):

I checked if there was mistakes in the question on my online school and I dont see any mistakes.

jhonyy9 (jhonyy9):

8/18 = aprox 0,4 from these choices 0,4 you will get fro 2/5 so this mean that your right answer will be C.

jhonyy9 (jhonyy9):

hope you understand it how you can solving same exercises

OpenStudy (denonakavro):

Jhony's answer is the best answer for this case, but the question asks for them to be proportional, 8/18 = 0.444444444, but 2/5 is 0.4, so they are the closest out of the other choices, but they are not proportional